Circular No. 6431

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

CYGNUS X-1 L. Piro, Istituto di Astrofisica Spaziale (IAS), CNR, Frascati; R. C.

Butler, Agenzia Spaziale Italiana; G. Boella, Istituto di Fisica Cosmica e Tecnologie Relative (IFCTR), Milan; L. Scarsi, Istituto di Fisica Cosmica e Applicazioni dell’Informatica (IFCAI), Palermo; J. A. M. Bleeker, Space Research Organization of the Netherlands (SRON), Utrecht; A. Parmar, Space Science Department, European Space Agency, Noordwijk; L. Chiappetti, IFCTR; B. Sacco, S. Giarrusso, and A. Santangelo, IFCAI; F. Frontera and D. Dal Fiume, Istituto Tecnologie e Studio delle Radiazioni Extraterrestri, Bologna; E. Costa, IAS; R. Jager and J. Heise, SRON; and P. Giommi, SAX Scientific Data Center, Rome, report on behalf of the SAX Team: “BeppoSAX successfully performed its first scientific observation by pointing at Cyg X-1 from June 22 to 25. All of the instruments performed nominally and showed that the source remains in its high state with a soft spectrum in the energy band 0.1–10 keV with an intensity of ∼ 1 Crab. Below 10 keV, the spectral shape is similar to that observed by ASCA (IAUC 6415). A hard component with a flux of ∼ 0.3 Crab is visible in the energy range 10–300 keV. The source intensity is variable with a peak amplitude of ∼ 40 percent at low and high energies. Typical intensities observed by the different instruments are: LECS (0.1–10 keV), 350 counts/s; MECS (1– 10 keV), 430 counts/s; HPGSPC (4–120 keV), 140 counts/s; PDS (15-300 keV), 65 counts/s; WFC (2-30 keV), 170 counts/s. Corresponding fluxes, in units of 10^{−8} erg cm^{−2} s^{−1}, for various ranges are: 0.1–2 keV, 1.3; 2–10 keV, 1.1; 4–34 keV, 1.1; 15–300 keV, 0.9; 2–30 keV, 1.6.”

DN GEMINORUM A. Retter and E. M. Leibowitz, Wise Observatory, Tel Aviv University,

write: “We performed CCD R photometry of the classical nova DN Gem during 14 nights in 1992 and 1993, and I photometry during 17 nights in 1995–1996. The light curve is modulated with the period 0.12785 ± 0.00005 day, with a peak-to-peak amplitude of 0.06 mag.”

COMET C/1996 N1 (BREWINGTON) Visual m_{1} and coma-diameter estimates: July 6.21 UT, 9.2, 4′.5 (C. S.

Morris, Pine Mountain Club, CA, 0.26-m reflector); 6.92, 9.1, 6′ (J. M. San Juan, Ciudad Real, Spain, 0.44-m reflector); 7.40, 10.0, 5′ (P. Camilleri, Cobram, Victoria, Australia, 20×80 binoculars); 8.18, 9.6, 3′.5 (A. Hale, Cloudcroft, NM, 0.41-m reflector).

1996 July 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green